OpenStudy (scienceteen8):

Sue has 7/8 of her book left to read. If she reads another 2/7 of the book, how much of the book will she have left? A.33/56 B.1/5 C.5/8 D.23/40

OpenStudy (jackofalltradez):

Do you have any ideas of the answer or how to find the answer?

OpenStudy (jackofalltradez):

@scienceteen8

OpenStudy (scienceteen8):

no

OpenStudy (jackofalltradez):

Ok, well, the first step is to find the common denominator of both fractions Do you know how to do that?

OpenStudy (scienceteen8):

multiply the numerator and denominator?

OpenStudy (jackofalltradez):

That would work in this problem But keep in mind sometimes there is a lower common denominator Anyways You would multiply both denominators to find the common denominator

OpenStudy (scienceteen8):

ok

OpenStudy (scienceteen8):

is the answer a

OpenStudy (jackofalltradez):

Why do you believe the answer is A?

OpenStudy (anonymous):

yup it's a

OpenStudy (anonymous):

(7/8)*(7/7)=(49/56) (2/7)*(8/8)= (16/56) So (49/56)-(16/56)= 33/56

OpenStudy (jackofalltradez):

@raventayler13 You would be absolutely correct however, I was trying to get @scienceteen8 to complete the problem so that they know how to do it themselves

OpenStudy (jackofalltradez):

Hence, they would learn how to do it on their own

OpenStudy (scienceteen8):

i did complete it lol i said it was a befor @raventayler13 completed it lol

OpenStudy (anonymous):

sorry @JackofallTradez i was just trying to help

OpenStudy (jackofalltradez):

And I was asking you why you thought it was A I didn't see any of your work, so I couldn't know if you had solved it or not

OpenStudy (jackofalltradez):

@raventayler13 I know, it's ok I was sort of tempted of just working out the problem and leaving the work to show my answer because this was such an easy problem for me

OpenStudy (scienceteen8):

oh ok i thought it as a because 7*8 is 56 and no i did not look at what @raventayler13 said bc i was trying to see if i got it right

OpenStudy (jackofalltradez):

@scienceteen8 Ah, that's why I thought you asked if A was the answer You can't always find the common denominator and assume that the answer with the common denominator is the correct answer, because sometimes the fraction simplifies

OpenStudy (jackofalltradez):

However, in this problem, it would be correct

OpenStudy (scienceteen8):

ok thank you @JackofallTradez

OpenStudy (jackofalltradez):

No problem, @scienceteen8 just here to help and explain math to others :P

OpenStudy (jackofalltradez):

And thank you for your input, @raventayler13

OpenStudy (anonymous):

Welcome @JackofallTradez i try

OpenStudy (jackofalltradez):

No problem I do believe we all try :P Nothing wrong with that

OpenStudy (anonymous):

lol can i have a medal?

OpenStudy (jackofalltradez):

Sure I knew I forgot something XD

OpenStudy (anonymous):

lol thank you @JackofallTradez

OpenStudy (jackofalltradez):

No problem You guys all have a nice day

OpenStudy (scienceteen8):

you to

OpenStudy (anonymous):

yup bye @scienceteen8 & @JackofallTradez